# Heat current and fluctuations between a dissipative qubit and a monitor under continuous measurement and feedback

### v2

@@ Abstract
Continuous quantum measurement and feedback induce heat exchange between a dissipative qubit and a monitor even in the steady state, as a measurement backaction. Using the Lindblad equation, we identified the maximum and minimum values of the steady-state heat current as the measurement and feedback states vary, and we demonstrate the qubit cooling induced by these processes. Turning our attention to quantum trajectories under continuous measurement and feedback, we observe that the heat current fluctuates around the steady-state values. We reveal that the fluctuations are strongly influenced by the measurement backaction, distinguishing them from the standard Poisson noise typically observed in electronic circuits. Our results offer potential application in the development of quantum refrigerators controlled by continuous measurement and feedback, and provide deep insight into quantum thermodynamics from the perspective of fluctuation.
@@ 1 Introduction
Quantum measurement and feedback induce unique non-unitary dynamics in a measured quantum system [1, 2, 3]. This measurement backaction has advanced the fields of quantum thermodynamics [4, 5, 6] and the application of quantum thermal machines [7, 8, 9, 10, 11, 12, 13]. The heat change in the measured quantum system due to the measurement and the feedback is a key concept in quantum energetics, e.g., work or heat extraction by operating measurement and feedback [14, 15]. Recent advances in quantum technology have allowed us to observe the heat exchange between the measured quantum system and the monitor under the measurement, which does not commute with the system Hamiltonian, in superconducting circuits [16]. This experimental progress motivates further theoretical research on the heat exchange by the non-commuting measurements. While the projective quantum measurement and feedback have been extensively studied so far [17, 18, 19, 20, 21, 22, 23, 24], recently, the continuous quantum measurement and feedback attract attention from fundamental aspects of quantum thermodynamics, such as the quantum fluctuation theorem [25] and quantum thermodynamics uncertainty relations (QTURs) [26, 27], as well as quantum thermal machines [28, 29, 30]. Most studies on quantum measurement have been conducted in the context of ideal systems without dissipation. It is crucial to take dissipation into account not only because of its ubiquity in nature but also due to its qualitative effects on heat exchange. This consideration enables us to address steady-state properties of heat exchange under continuous quantum measurement and feedback, as measurement and feedback alone induce no heat exchange in the steady-state limit.
@@
Continuous quantum measurement and feedback process is inherently stochastic, as the quantum state probabilistically jumps to a measured state. Consequently, the heat current induced by continuous quantum measurement and feedback fluctuates around the ensemble-averaged heat current with respect to the measurement outcomes. Current fluctuations in quantum transport have been studied in the mesoscopic physics [31, 32]. For instance, for non-interacting electrons in a one-dimensional nanowire with weak transmission, the variance of the electronic current, which characterizes current fluctuations, is proportional to the mean electronic current. This is reflected by the fact that free electrons transferring through the nanowire obey the Poisson statistics. Observing current fluctuations provides deeper insights into the microscopic scattering process of carriers. Now, current fluctuations have been studied theoretically and experimentally in strongly correlated systems, such as the fractional quantum Hall effect [33, 34], the Kondo effect [35, 36], and superconductivity [37, 38, 39]. Moreover, since the statistical properties lie behind the current noise, it is closely related to the fluctuation theorem [40, 41].
@@
In this paper, we investigate the heat exchange between a monitor and a qubit with dissipation by connecting the qubit to bosonic environments, under continuous quantum measurement and feedback. In the first part, we examine the steady-state heat current using the Lindblad equation and derive its upper and lower bounds. Reference [42] has demonstrated that, in the absence of feedback, the heat current does not flow out of the dissipative qubit to the monitor under continuous quantum measurement. In this work, we incorporate a feedback sequence following quantum measurement to create and stabilize desired quantum states that cannot be achieved through measurement alone. As a result, we demonstrate that the monitor can extract heat from the dissipative qubit. This cooling effect could have potential applications in measurement-based quantum refrigerators. The second part focuses on the fluctuations in the heat current by tracing individual quantum trajectories. We demonstrate that the power spectrum consists of two distinct processes unique to continuous quantum measurement: Poisson noise and backaction noise, associated with quantum jumps and measurement backaction dynamics, respectively. Notably, we observe that the Fano factor, which is the ratio of total noise to Poisson noise, remains below unity across a wide range of parameters. The perspective from fluctuations of the continuous quantum measurement and feedback can lead to the further understating of thermodynamics in quantum dissipative systems.
@@
This paper is organized as follows. In the next section, we introduce the theoretical model used in this study and describe quantum dynamics with both dissipation and continuous measurement and feedback. In Sec. 3, we calculate the steady-state heat current between the dissipative qubit and the monitor and obtain its upper and lower bounds. In Sec. 4, we address the fluctuations of the heat current and analyze the noise power spectra. Finally, the summary is drawn in Sec. 5. Throughout this paper, We set $\hbar=k_{B}=1$.
@@ 2 Model
Figure 1: Monitored qubit is coupled to the heat baths. The qubit state is continuously measured with the strength $\gamma$, and affected by the measurement as a backaction.
@@ 2.2 Measurement and feedback
$\mathcal{D}_{\rm B}[\rho(t)]{\color[rgb]{0,0,0}=\sum_{r}\mathcal{D}_{{\rm B},r}[\rho(t)]}=\sum_{s=\pm}\Gamma_{s}\left(\sigma_{\bar{s}}\rho(t)\sigma_{s}-\frac{1}{2}\left\{\sigma_{s}\sigma_{\bar{s}},\rho(t)\right\}\right),$ (5)
with $\bar{s}=-s$. Here, $\Gamma_{+}=\sum_{r}(\pi/2)I_{r}(\Delta)[1+n_{r}(\Delta)]$ and $\Gamma_{-}=\sum_{r}(\pi/2)I_{r}(\Delta)n_{r}(\Delta)$ are the total emission and absorption rates of multiple heat baths, respectively, where $I_{r}(\omega)=\sum_{k}\lambda_{rk}^{2}\delta(\omega-\omega_{rk})$ is the spectral density for the heat bath $r$. In this work, we assume the Ohmic heat bath, $I_{r}(\omega)=2\alpha_{r}\omega e^{-\omega/\omega_{\rm c}}$ with the dimensionless coupling strength $\alpha_{r}$ and the cutoff frequency $\omega_{\rm c}$ [43, 44]. Note that, to justify the Born-Markov approximation, the total emission and absorption rates $\Gamma_{\pm}$ must be much smaller compared with the qubit energy $\Delta$ [45]. As long as this condition is justified, it does not matter if the heat baths have different temperatures.
3 Steady-state heat current
We are interested in the heat flow from the monitor into the qubit [29, 42],
@@ 3 Steady-state heat current
where ${\rm tr}_{0}[\cdots]$ denotes the trace about the degrees of freedom of the qubit. The positive (negative) heat current indicates the heating (cooling) of the qubit by the quantum measurement and feedback.
@@ 3.1 Minimum and maximum
At the steady-state limit, we can evaluate the steady-state heat current by solving the Lindblad equation (4) for $\dot{\rho}(t)=0$. For the weak coupling case ($\gamma,~{}\Gamma_{\pm}\ll\Delta$), the steady-state heat current is approximated as
@@
where $\gamma_{\pm}=\Gamma_{+}\pm\Gamma_{-}$, thus, $0<\gamma_{-}<\gamma_{+}$. Note that the dependence of the steady-state heat current (7) on $\phi_{m}$ and $\phi_{n}$ is a minor correction for the weak coupling case. Thus, the measurement and feedback states are characterized by $\theta_{m}$ and $\theta_{n}$, respectively. The steady-state heat current satisfies the following inequality,
@@
The maximum and minimum heat currents $J_{{\rm max}/{\rm min}}=\pm\gamma\Delta\Gamma_{\pm}/(\gamma_{+}+\gamma)$ occur at $(\theta_{m},\theta_{n})=(0,\pi)$ and $(\pi,0)$, respectively. This inequality indicates that the qubit can be both heated and cooled by the monitoring and feedback.
For $(\theta_{m},\theta_{n})=(\pi,0)$, the measurement is to the excited state and then feedback is to the ground state. In this process, the qubit energy decreases by $\Delta$, cooling the qubit. The steady-state heat current takes minimum $J_{\rm min}=-\gamma\rho_{ee}(\bra{e}H_{0}\ket{e}-\bra{g}H_{0}\ket{g})=-\gamma\Delta\Gamma_{-}/(\gamma_{+}+\gamma)$, where $\rho_{ij}=\bra{i}\rho\ket{j}$. On the other side, for $(\theta_{m},\theta_{n})=(0,\pi)$, the measurement projects the system in the ground state and the feedback in the excited state. For this case, the qubit energy is increased by the measurement and the feedback, changing the qubit energy by $\Delta$. The steady-state heat current is maximum and written as $J_{\rm max}=-\gamma\rho_{gg}(\bra{g}H_{0}\ket{g}-\bra{e}H_{0}\ket{e})=\gamma\Delta\Gamma_{+}/(\gamma_{+}+\gamma)$.
Figure 2: Steady-state heat current as a function of the measurement state $\theta_{m}$ and the feedback state $\theta_{n}$ for $\Gamma_{+}/\Delta=0.1$, $\Gamma_{-}/\Delta=0.05$, and $\gamma/\Delta=0.1$. These parameters correspond to $\alpha_{\rm eff}\approx 0.0191$ and $T_{\rm eff}/\Delta\approx 1.44$. For the measurement-only case ($\theta_{m}=\theta_{n}$), the heat current never takes negative values.
Figure 2 shows the steady-state heat current for $\Gamma_{+}/\Delta=0.1$, $\Gamma_{-}/\Delta=0.05$, and $\gamma/\Delta=0.1$, which corresponds to the effective coupling strength $\alpha_{\rm eff}\approx 0.0191$ and the effective temperature $T_{\rm eff}/\Delta\approx 1.44$ when the effects of multiple heat baths are represented by an effective single heat bath. The steady-state heat current reaches maximum at $(\theta_{m},\theta_{n})=(0,\pi)$ and minimum at $(\theta_{m},\theta_{n})=(\pi,0)$, as expected, and its sign changes between them. The heat current does not flow at the edges $(\theta_{m},\theta_{n})=(0,0)$ and $(\pi,\pi)$ and on the convex curve towards the measurement state $\theta_{m}$ connecting the edges. The convex curve indicates that the steady-state heat current never takes negative ($J\geq 0$) for the measurement only case, $\theta_{n}=\theta_{m}$ [42].
Note that the heat current can exceed the steady-state bounds. When the qubit is prepared in the excited (ground) state for measurement to the excited (ground) state and feedback to the ground (excited) state, the heat current reaches its minimum (maximum) value, $\mp\gamma\Delta$.
@@ 3.2 Symmetry at $\theta_{m}+\theta_{n}=\pi$
The steady-state heat current is symmetric with $\theta_{m}+\theta_{n}=\pi$, corresponding to the feedback of the measurement state to the opposite side of the Bloch sphere, in the weak coupling case. On the axis of the mirror symmetry, $\theta_{m}+\theta_{n}=\pi$, the steady-state heat current reads
@@
This monotonically decreases from the maximum to the minimum heat current as a function of $\theta_{m}$, and its sign changes at $\theta_{m}=\pi-\cos^{-1}[\tanh(\Delta/4T)]$ ($>\pi/2$). As the temperature decreases, the positive heat current area is extended. This asymmetry between the positive and negative heat currents is controlled by the temperature of the heat baths coupled with the qubit. For $\theta_{m}=|\epsilon|$ and $\theta_{m}=\pi-|\epsilon|$, where the steady-state heat current goes to maximum and minimum at $|\epsilon|\to 0$, respectively, the steady-state heat current deviates quadratically from the maximum and minimum heat currents in the same way for $|\epsilon|\ll 1$,
@@
3.3 Zero heat current
Finally, we consider the condition of the zero heat current ($J=0$). At the edges, $(\theta_{m},\theta_{n})=(0,0)$ and $(\pi,\pi)$, the measurement is commuting with the qubit Hamiltonian. The commuting measurement does not disturb the qubit state, and then the steady-state heat current does not flow. When the feedback process is present, |$\ket{m}\neq\ket{n}$, the qubit state is disturbed by the measurement and feedback process. Near the lower-left edge, where $\theta_{m},\theta_{n}\ll 1$, the heat current vanishes when $\theta_{n}=e^{-\Delta/(2T)}\theta_{m}$. Conversely, near the upper-right edge, where $\pi-\theta_{m},\pi-\theta_{n}\ll 1$, heat flow ceases when $\pi-\theta_{n}=e^{+\Delta/(2T)}(\pi-\theta_{m})$. As the temperature decreases, the curve in which the steady-state heat current vanishes deviates more from the diagonal line ($\theta_{n}=\theta_{m}$), and then the positive heat current area is extended.

### v3

# Energy exchange and fluctuations between a dissipative qubit and a monitor under continuous measurement and feedback

@@ Abstract
Continuous quantum measurement and feedback induce energy exchange between a dissipative qubit and a monitor even in the steady state, as a measurement backaction. Using the Lindblad equation, we identified the maximum and minimum values of the steady-state energy flow as the measurement and feedback states vary, and we demonstrate the qubit cooling induced by these processes. Turning our attention to quantum trajectories under continuous measurement and feedback, we observe that the energy flow fluctuates around the steady-state values. We reveal that the fluctuations are strongly influenced by the measurement backaction, distinguishing them from the standard Poisson noise typically observed in electronic circuits. Our results offer potential application in the development of quantum refrigerators controlled by continuous measurement and feedback, and provide deep insight into quantum thermodynamics from the perspective of fluctuation.
@@ 1 Introduction
Quantum measurement and feedback induce unique non-unitary dynamics in a measured quantum system [1, 2, 3]. This measurement backaction has advanced the fields of quantum thermodynamics [4, 5, 6] and the application of quantum thermal machines [7, 8, 9, 10, 11, 12, 13]. The heat change in the measured quantum system due to the measurement and the feedback is a key concept in quantum energetics, e.g., work or heat extraction by operating measurement and feedback [14, 15]. Recent advances in quantum technology have allowed us to observe the energy exchange between the measured quantum system and the monitor under the measurement, which does not commute with the system Hamiltonian, in superconducting circuits [16]. This experimental progress motivates further theoretical research on the energy exchange by the non-commuting measurements. While the projective quantum measurement and feedback have been extensively studied so far [17, 18, 19, 20, 21, 22, 23, 24], recently, the continuous quantum measurement and feedback attract attention from fundamental aspects of quantum thermodynamics, such as the quantum fluctuation theorem [25] and quantum thermodynamics uncertainty relations (QTURs) [26, 27], as well as quantum thermal machines [28, 29, 30]. Most studies on quantum measurement have been conducted in the context of ideal systems without dissipation. It is crucial to take dissipation into account not only because of its ubiquity in nature but also due to its qualitative effects on energy exchange. This consideration enables us to address steady-state properties of energy flow under continuous quantum measurement and feedback, as measurement and feedback alone induce no energy exchange in the steady-state limit.
@@
Continuous quantum measurement and feedback process is inherently stochastic, as the quantum state probabilistically jumps to a measured state. Consequently, the energy exchange induced by continuous quantum measurement and feedback fluctuates around the ensemble-averaged energy flow with respect to the measurement outcomes. Current fluctuations in quantum transport have been studied in the mesoscopic physics [31, 32]. For instance, for non-interacting electrons in a one-dimensional nanowire with weak transmission, the variance of the electronic current, which characterizes current fluctuations, is proportional to the mean electronic current. This is reflected by the fact that free electrons transferring through the nanowire obey the Poisson statistics. Observing current fluctuations provides deeper insights into the microscopic scattering process of carriers. Now, current fluctuations have been studied theoretically and experimentally in strongly correlated systems, such as the fractional quantum Hall effect [33, 34], the Kondo effect [35, 36], and superconductivity [37, 38, 39]. Moreover, since the statistical properties lie behind the current noise, it is closely related to the fluctuation theorem [40, 41].
@@
In this paper, we investigate the energy exchange between a monitor and a qubit with dissipation by connecting the qubit to bosonic environments, under continuous quantum measurement and feedback. In the first part, we examine the steady-state energy flow using the Lindblad equation and derive its minimum and maximum values by varying the measurement and feedback states. Reference [42] has demonstrated that, in the absence of feedback, the energy exchange does not occur from the dissipative qubit to the monitor under continuous quantum measurement. In this work, we incorporate a feedback sequence following quantum measurement to create and stabilize desired quantum states that cannot be achieved through measurement alone. As a result, we demonstrate that the monitor can extract energy from the dissipative qubit. This cooling effect could have potential applications in measurement-based quantum refrigerators. The second part focuses on the fluctuations in the energy flow by tracing individual quantum trajectories. We demonstrate that the power spectrum consists of two distinct processes unique to continuous quantum measurement: Poisson noise and backaction noise, associated with quantum jumps and measurement backaction dynamics, respectively. Notably, we observe that the Fano factor, which is the ratio of total noise to Poisson noise, remains below unity across a wide range of parameters. The perspective from fluctuations of the continuous quantum measurement and feedback can lead to the further understating of thermodynamics in quantum dissipative systems.
@@
This paper is organized as follows. In the next section, we introduce the theoretical model used in this study and describe quantum dynamics with both dissipation and continuous measurement and feedback. In Sec. 3, we calculate the steady-state energy flow between the dissipative qubit and the monitor and obtain its minimum and maximum values. In Sec. 4, we address the fluctuations of the energy exchange and analyze the noise power spectra. Finally, the summary is drawn in Sec. 5. Throughout this paper, We set $\hbar=k_{B}=1$.
@@ 2 Model
Figure 1: Monitored qubit is coupled to the heat baths. The qubit state is continuously measured with the strength $\gamma$, and affected by the measurement as a backaction and feedback.
@@ 2.2 Measurement and feedback
$\mathcal{D}_{\rm B}[\rho(t)]=\sum_{r}\mathcal{D}_{{\rm B},r}[\rho(t)]=\sum_{s=\pm}\Gamma_{s}\left(\sigma_{\bar{s}}\rho(t)\sigma_{s}-\frac{1}{2}\left\{\sigma_{s}\sigma_{\bar{s}},\rho(t)\right\}\right),$ (5)
with $\bar{s}=-s$. Here, $\Gamma_{+}=\sum_{r}(\pi/2)I_{r}(\Delta)[1+n_{r}(\Delta)]$ and $\Gamma_{-}=\sum_{r}(\pi/2)I_{r}(\Delta)n_{r}(\Delta)$ are the total emission and absorption rates of multiple heat baths, respectively, where $I_{r}(\omega)=\sum_{k}\lambda_{rk}^{2}\delta(\omega-\omega_{rk})$ is the spectral density for the heat bath $r$. In this work, we assume the Ohmic heat bath, $I_{r}(\omega)=2\alpha_{r}\omega e^{-\omega/\omega_{\rm c}}$ with the dimensionless coupling strength $\alpha_{r}$ and the cutoff frequency $\omega_{\rm c}$ [43, 44]. Note that, to justify the Born-Markov approximation, the total emission and absorption rates $\Gamma_{\pm}$ must be much smaller compared with the qubit energy $\Delta$ [45]. As long as this condition is met, it does not matter if the heat baths have different temperatures.
3 Steady-state energy flow
We are interested in the energy flow from the monitor into the qubit [29, 42],
@@ 3 Steady-state energy flow
where ${\rm tr}_{0}[\cdots]$ denotes the trace about the degrees of freedom of the qubit. The positive (negative) energy flow indicates the heating (cooling) of the qubit by the quantum measurement and feedback.
@@ 3.1 Minimum and maximum
At the steady-state limit, we can evaluate the steady-state energy flow by solving the Lindblad equation (4) for $\dot{\rho}(t)=0$. For the weak coupling case ($\gamma,~{}\Gamma_{\pm}\ll\Delta$), the steady-state energy flow is approximated as
@@
where $\gamma_{\pm}=\Gamma_{+}\pm\Gamma_{-}$, thus, $0<\gamma_{-}<\gamma_{+}$. Note that the dependence of the steady-state energy flow (7) on $\phi_{m}$ and $\phi_{n}$ is a minor correction for the weak coupling case. Thus, the measurement and feedback states are characterized by $\theta_{m}$ and $\theta_{n}$, respectively. The steady-state energy flow satisfies the following inequality,
@@
The maximum and minimum energy flows $J_{{\rm max}/{\rm min}}=\pm\gamma\Delta\Gamma_{\pm}/(\gamma_{+}+\gamma)$ occur at $(\theta_{m},\theta_{n})=(0,\pi)$ and $(\pi,0)$, respectively. This inequality indicates that the qubit can be both heated and cooled by the monitoring and feedback.
For $(\theta_{m},\theta_{n})=(\pi,0)$, the measurement is to the excited state and then feedback is to the ground state. In this process, the qubit energy decreases by $\Delta$, cooling the qubit. The steady-state energy flow takes minimum $J_{\rm min}=-\gamma\rho_{ee}(\bra{e}H_{0}\ket{e}-\bra{g}H_{0}\ket{g})=-\gamma\Delta\Gamma_{-}/(\gamma_{+}+\gamma)$, where $\rho_{ij}=\bra{i}\rho\ket{j}$. On the other side, for $(\theta_{m},\theta_{n})=(0,\pi)$, the measurement projects the system in the ground state and the feedback in the excited state. For this case, the qubit energy is increased by the measurement and the feedback, changing the qubit energy by $\Delta$. The steady-state energy flow is maximum and written as $J_{\rm max}=-\gamma\rho_{gg}(\bra{g}H_{0}\ket{g}-\bra{e}H_{0}\ket{e})=\gamma\Delta\Gamma_{+}/(\gamma_{+}+\gamma)$.
Figure 2: Steady-state energy flow as a function of the measurement state $\theta_{m}$ and the feedback state $\theta_{n}$ for $\Gamma_{+}/\Delta=0.1$, $\Gamma_{-}/\Delta=0.05$, and $\gamma/\Delta=0.1$. These parameters correspond to $\alpha_{\rm eff}\approx 0.0191$ and $T_{\rm eff}/\Delta\approx 1.44$. For the measurement-only case ($\theta_{m}=\theta_{n}$), the energy flow never takes negative values.
Figure 2 shows the steady-state energy flow for $\Gamma_{+}/\Delta=0.1$, $\Gamma_{-}/\Delta=0.05$, and $\gamma/\Delta=0.1$, which corresponds to the effective coupling strength $\alpha_{\rm eff}\approx 0.0191$ and the effective temperature $T_{\rm eff}/\Delta\approx 1.44$ when the effects of multiple heat baths are represented by an effective single heat bath. The steady-state energy flow reaches maximum at $(\theta_{m},\theta_{n})=(0,\pi)$ and minimum at $(\theta_{m},\theta_{n})=(\pi,0)$, as expected, and its sign changes between them. The energy exchange does not occur at the edges $(\theta_{m},\theta_{n})=(0,0)$ and $(\pi,\pi)$ and on the convex curve towards the measurement state $\theta_{m}$ connecting the edges. The convex curve indicates that the steady-state energy flow never takes negative ($J\geq 0$) for the measurement only case, $\theta_{n}=\theta_{m}$ [42].
Note that the energy flow can exceed the minimum and maximum values of the steady-state energy flow. When the qubit is prepared in the excited (ground) state for measurement to the excited (ground) state and feedback to the ground (excited) state, the energy flow reaches its minimum (maximum) value, $\mp\gamma\Delta$.
@@ 3.2 Symmetry at $\theta_{m}+\theta_{n}=\pi$
The steady-state energy flow is symmetric with $\theta_{m}+\theta_{n}=\pi$, corresponding to the feedback of the measurement state to the opposite side of the Bloch sphere, in the weak coupling case. On the axis of the mirror symmetry, $\theta_{m}+\theta_{n}=\pi$, the steady-state energy flow reads
@@
This monotonically decreases from the maximum to the minimum energy flow as a function of $\theta_{m}$, and its sign changes at $\theta_{m}=\pi-\cos^{-1}[\tanh(\Delta/4T)]$ ($>\pi/2$). As the temperature decreases, the positive energy flow area is extended. This asymmetry between the positive and negative energy flows is controlled by the temperature of the heat baths coupled with the qubit. For $\theta_{m}=|\epsilon|$ and $\theta_{m}=\pi-|\epsilon|$, where the steady-state energy flow goes to maximum and minimum at $|\epsilon|\to 0$, respectively, the steady-state energy flow deviates quadratically from the maximum and minimum energy flows in the same way for $|\epsilon|\ll 1$,
@@
3.3 Zero energy exchange
Finally, we consider the condition of the zero energy exchange ($J=0$). At the edges, $(\theta_{m},\theta_{n})=(0,0)$ and $(\pi,\pi)$, the measurement is commuting with the qubit Hamiltonian. The commuting measurement does not disturb the qubit state, and then the steady-state energy flow does not flow. When the feedback process is present, |$\ket{m}\neq\ket{n}$, the qubit state is disturbed by the measurement and feedback process. Near the lower-left edge, where $\theta_{m},\theta_{n}\ll 1$, the energy flow vanishes when $\theta_{n}=e^{-\Delta/(2T)}\theta_{m}$. Conversely, near the upper-right edge, where $\pi-\theta_{m},\pi-\theta_{n}\ll 1$, energy flow ceases when $\pi-\theta_{n}=e^{+\Delta/(2T)}(\pi-\theta_{m})$. As the temperature decreases, the curve in which the steady-state energy flow vanishes deviates more from the diagonal line ($\theta_{n}=\theta_{m}$), and then the positive energy flow area is extended.